OpenStudy (anonymous):

Solve x2 + 10x = -21

OpenStudy (anonymous):

What do you have to solve there? Solve for x?

OpenStudy (anonymous):

I would solve using the quadratic formula

OpenStudy (anonymous):

Guessing method?

OpenStudy (anonymous):

so \[x=((-b)\pm \sqrt{b^2-4ac)})/2a\]

OpenStudy (anonymous):

where ax^2 + bx = -c are your a, b, and c values

OpenStudy (anonymous):

so for your problem. 1 = a, 10 = b and 21 = c (if you add the 21 to both sides and make it positive)

OpenStudy (anonymous):

so plug those numbers into your equation

OpenStudy (anonymous):

so plug those numbers into your equation\[x= (-(10) \pm \sqrt{(10)^2-4(1)(21)}/2(1)\]

OpenStudy (anonymous):

and solve from there

OpenStudy (anonymous):

@Katerrah did you get your two answers?

OpenStudy (anonymous):

Can you be clear on what you're trying to solve for, hun?

OpenStudy (aum):

x^2 + 10x = -21 x^2 + 10x + 21 = 0 x^2 + 3x + 7x + 21 = 0 x(x+3) + 7(x+3) = 0 (x+3)(x+7) = 0 x = -3 or x = -7

geerky42 (geerky42):

@TechnoSoul Use `\dfrac{}{}` next time. it makes fraction more readable example: `\dfrac{1}{2}` display \(\dfrac{1}{2}\)

OpenStudy (anonymous):

x=-7, -3 actually, I don't think there should be an or?